LARENZIE:

Solve the following equation -2(-3x + 6x) = 30 A x=5 B x=6 C x=-6 D x=-5

LARENZIE:

help pls

SmokeyBrown:

To solve this equation, let's try to get "x" by itself. I'll show you how to do it step by step, so we can get the answer and you can understand the method. -2(-3x + 6x) = 30 We can start by multiplying the -2 outside the parentheses by everything inside the parentheses. This is called distribution. The result of this is: 6x - 12x = 30 Next, we can combine 6x and -12 x: -6x = 30 Finally, we will do a simple division, dividing both sides of the equation by negative 6. What we get at the end is: x = ___ Have you followed along so far, and can you fill in that last blank to get your answer?

LARENZIE:

thx love

LARENZIE:

-6

SmokeyBrown:

If x is equal to -6, then we can try replacing x with -6: We had -6x = 30 Which becomes -6 * (-6) = 30 This is not true; -6*(-6) is actually 36, not 30. Please feel free to try again; you are pretty close

LARENZIE:

-5

SmokeyBrown:

There you go! Well done. -6 * (-5) is indeed equal to 30. Looks like you've found your answer :)

LARENZIE:

Solve the following equation 6(4x - x) = -36 A x=2 B x=-4 C x=-2 D x=4

SmokeyBrown:

Normally it would be better to make a new post for a separate question, but this question is very closely related to the one we already did, so I don't think it's a big deal. Since the steps you'd take to solve this problem are very very similar to the previous one, I will just refer you to the method we used above and trust you to do the calculations yourself this time. It is essentially the same question with different numbers. Remember, to solve for "x" you want to get "x" by itself on one side of the equation. As always, if there is anything else you need, please do not hesitate to ask.